# DRL-Based Resource Allocation for Motion Blur Resistant Federated Self-Supervised Learning in IoV

### v1

@@ I Introduction
Internet of Vehicles (IoV) is a technology that connects vehicles, road and transportation facilities to Internet, enabling intelligent interaction and information sharing among them. Advances in IoV have made many practical applications possible, e.g., automatic navigation, meteorological information and road condition monitoring, which provide convenience to drivers and self-driving system, reduce the probability of accidents and improve the efficiency of the transportation system [1]-[3]. These applications need robust models, which require a lot of data for training. In IoV, vehicles can continuously capture the fresh data along the moving road, including road conditions and environmental states through onboard sensors, where most of the data are captured in the form of images [4]-[6]. Moreover, vehicles also can store the image data locally and deal with them to realize recognition and classification, which provides essential information for self-driving and driver assistance systems to aid drivers in perceiving and understanding surroundings.
@@
However, the locally stored image data, denoted as local data, is usually private for vehicles, and vehicles are reluctant to share the data with each other to protect privacy. Since vehicles come from all directions, the local data they collect vary greatly. If the model, denoted as local model, is trained locally based on the local data, it may converge to a local optimum [7]-[9]. In contrast to distributed training, Federated Learning (FL) performs a centered training and enables a Base Station (BS) to obtain a global model through aggregating local models from vehicles without sharing the local data, thereby providing a privacy-preserving solution and reducing the differences between local models [10][11]. For local training, the traditional supervised learning requires the image data with labels [12]-[14], but data labeling requires significant manual effort. Especially in IoV, a large amount of unlabeled data is generated daily, further increasing the cost of manual labeling. The Self-Supervised Learning (SSL) leverages pretext to provide training based on the inherent properties of data, thus it can learn feature representations from the data itself without the need for manually annotated labels. The traditional SSL methods obtain a better training performance through generating and storing abundant negative samples, which occupies a lot of computing resource and storage space. Thus, traditional methods are not suitable for the resource and space limited IoV. The famous SSL method Momentum Contrast (MoCo) uses $k$ values to form a dictionary and stand for the reference of negative samples, which can save computing resource and storage space [15][16]. However, employing MoCo in FSSL for local training will upload $k$ values to create a big dictionary in BS for aggregation, which may cause privacy leakage [17]-[19]. Moreover, the training performance of FSSL in IoV may be deteriorated due to the motion blur. Specifically, excessive vehicle velocity may result in insufficient exposure time of the camera sensor, causing image blur [20][21]. The blur level of images directly impacts the accuracy and convergence of the global model. To the best of our knowledge, there is no work considering the privacy leakage and motion blur for FSSL in IoV.
@@
The improved MoCo method Simplified Contrast (SimCo) uses dual temperature to control sample distribution to eliminate the usage of dictionary [22]-[24], thus it can greatly protect privacy and is suitable for the local training of FSSL in IoV. Considering the negative impact of motion blur on model aggregation, we also propose a motion Blur resistant FSSL, refer to as BFSSL.
@@
It is noteworthy that increasing Central Processing Unit (CPU) frequency of a vehicle results in more local training iterations while consuming more energy consumption. Higher transmission power of a vehicle reduces the transmission delay, while costing more energy. Simultaneously, low CPU frequency reduces the number of local iterations within the given processing period of time, thereby diminishing the accuracy of the global model [25][26]. Additionally, low transmission power increases data interference, resulting in higher rates of transmission failure. Thus, it is critical to find a resource allocation scheme that allocates resources (i.e., CPU frequency and transmission power of vehicles) to minimize energy consumption and delay for the BFSSL, while increasing the success rate of transmission in IoV. Since the resource allocation problem is usually non-convex in the complicated IoV, the traditional convex optimization methods are difficult to solve it. Deep Reinforcement Learning (DRL) is a good option to solve the non-convex problem. In this paper, we propose the DRL-BFSSL algorithm, where vehicles employs SimCo to perform locally training and sends local models to BS for blur-based aggregation. Meanwhile, BS allocates CPU frequency and transmission power to minimize the energy consumption and delay [^1] The source code has been released at: https://github.com/qiongwu86/DRL-BFSSL.
@@
  • In IoV, we utilize the level of motion blur as a crucial weighting factor on Independent and Identically Distributed (IID) and Non-IID datasets during model aggregation. Subsequently, by aggregating models uploaded by training vehicles associated with the blur level, we obtain a blur resistant global model.
  • We firstly formulate an optimization problem aiming to minimize the sum of energy consumption and delay. Subsequently, utilizing the Karush-Kuhn-Tucker (KKT) conditions to determine the optimal solution that jointly allocating bandwidth and CPU frequency. Finally, by the Soft Actor-Critic (SAC) algorithm, we identify the optimal allocation schedule for the CPU frequency and transmission power of each vehicle. Meanwhile, considering that low transmission power may lead to data transmission failures, we use the data error rate to set a minimum transmission power limit.
The rest of this paper is divided into several interrelated parts to comprehensively explore the discussed problem. In Section II, we will review prior work relevant to this study, providing background and context for subsequent discussions. In Section III, we describe the system model. In Section IV, we will formulate an optimization problem aiming to minimize the sum of energy consumption and delay and in Section V, we will introduce our proposed DRL-BFSSL algorithm. Section VI showcase and discuss the simulation results. Ultimately, we will summarize the paper in Section VII.
@@ II-A Federated SSL
In [27], Jahan et al. proposed a FSSL which combines FL and the SSL method (i.e., Simple framework for Contrastive Learning of Representations (SimCLR)), to identify Mpox from skin lesion images [28]. In [29], Yan et al. proposed a FSSL which combines FL and the SSL method (i.e., Variational inference with adversarial learning for end-to-end text-to-speech (Vits)), and utilizes medical images stored in each hospital to train models. In [30], Xiao et al. proposed a FSSL which integrates Generative Adversarial Networks (GAN) as a pretext for SSL and combines it with FL to support automatic traffic analysis and synthesis over a large number of heterogeneous datasets. In [31], Li et al. proposed a FSSL which integrates You Only Look Once (YOLO) as a pretext for SSL and combines it with FL to process object detection in power operation sites. These FSSL methods mentioned above require a large number of samples for training, which imposes high demands on the computing and storage capabilities of devices. However, the computing and storage resources of vehicles are limited in IoV, thus these methods are not suitable for IoV.
@@ II-B DRL-based resource allocation scheme in FL
Some works have proposed DRL-based resource allocation schemes in IoV. In [36], Zheng et al. proposed a resource allocation scheme based on DRL method to minimize the duration required for accomplishing tasks while respecting latency limitations in IoV. In [37], Hazarika et al. proposed a resource allocation scheme based on DRL method to allocate the power optimally in IoV. In [38], Guo et al. proposed a resource allocation scheme based on DRL method to minimize the cost of Road Side Unit (RSU) in IoV. However, the methods mentioned above have not considered the FL framework. There are also some works that consider the FL framework to address resource allocation problems based on DRL. In [39], He et al. combined FL and DRL method (i.e., Deep Deterministic Policy Gradient (DDPG)), to optimize both computation offloading and resource allocation. In [40], Zhang et al. combined FL and DRL method (i.e., DDPG), and proposed a dynamic computation offloading and resource allocation scheme to minimize the delay and energy consumption.
However, there is no work concerning the DRL-based resource allocation schemes for FSSL in IoV. Therefore, we further investigate a DRL-based resource allocation scheme to optimize the performance for the BFSSL in IoV, referred to as DRL-BFSSL.
@@ III System Model
In this section, we will introduce the system scenario and models. As shown in Fig. 1, we consider a scenario of a BS deployed at the center of the intersection and several vehicles driving in the coverage of the BS. When a vehicle arrives at the intersection, it turns left, turns right and drives straightly with probabilities $\varepsilon_{1}$, $\varepsilon_{2}$, $\varepsilon_{3}$, and $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}=1$, and then it keeps driving straightly. Vehicle velocity follows the truncated Gaussian distribution. Each vehicle is equipped with a camera to capture images during driving through the coverage area of the BS, and the vehicles coming from different directions capture different categories of images. $N$ training vehicles are randomly selected from the vehicles before crossing the intersection to ensure that vehicles have enough time for local training. Vehicle $n$, $n\in[1,N]$, is equipped with a camera and captures $Z$ image data with velocity $v_{0}^{n}$ before it enters the coverage area of the BS.
The entire process of DRL-BFSSL algorithm consists of $K^{\text{max}}$ episodes, each of which is divided into two task lines, one for BFSSL, where episode $k$ is divided into different rounds, and the other for DRL-based resource allocation, where episode $k$ is divided into different slots[41][42].
Specifically, as shown in Fig. 2, within each episode $k$, $k\in[1,K^{\text{max}}]$ , there are $S^{\text{max}}$ slots, and each round $r$ corresponds to one slot. The difference is that at the end of each episode $k$, when entering the next episode $k+1$, slot $t$ is reset to slot 00. In contrast, round $r$ is not reset and continues to accumulate until the whole process of DRL-BFSSL algorithm finished. Due to each episode containing $S^{\text{max}}$ slots, and round $r$ resets to 0 after the completion of the $K^{\text{max}}$ episodes, the relationship between slot $t$ in episode $k$ and round $r$ is as follows
@@
At the beginning of the DRL-BFSSL algorithm, the BS collects the initial positions of training vehicles and initializes a global model. For the task line of BFSSL, at the beginning of each round $r$, BS uses the model aggregated at the end of round $r-1$ as the global model for round $r$. Meanwhile, for the task line of resource allocation, at the beginning of each slot $t$, the BS receives the velocities uploaded by vehicles and estimate the positions based on the initial positions and received velocities, after that it adopts the DRL method to allocate resources for training vehicles including transmission power and CPU frequency. Then each training vehicle downloads the global model from the BS, along with the allocated transmission power and CPU frequency. After that each vehicle trains locally based on the global model to update its local model. In this process, each vehicle firstly randomly chooses a certain number of images from captured images as local data and subsequently conducts local training through iterations to update the local model. During the process of local training, the vehicles also capture new images for the training of next round. When the local training is finished, each vehicle adopts the downloaded transmission power to transmit its updated local model and velocity to BS. However, during the transmission process, varying signals transmitted by other vehicles may cause interference, leading to data errors that prevent the BS from receiving the uploaded local models. Finally, BS aggregates the received local models to obtain a new global model based on the motion blur. The above process is repeated in different episode to get a converged global model.
@@ III-A Mobility model
As the velocity of each vehicle $n$ follows the truncated Gaussian distribution to reflect real vehicle mobility, the velocities of different vehicles are IID. Let $v_{k,t}^{n}$ be the velocity of vehicle $n$, and $v^{\text{min}}$ and $v^{\text{max}}$ be the minimum and maximum velocity of vehicles, respectively. Thus the probability density function of $v_{k,t}^{n}$ is calculated as [43]-[45]
@@ III-B Transmission model
During the transmission process, the energy consumption of vehicle $n$ can be calculated as [46][47]
@@
where $c_{k,t}^{n}$ is the transmission rate of vehicle $n$, which can be calculated according to Shannon’s theorem [48]-[50], i.e.,
@@
where $g_{n}$ represents the small-scale fast fading power, which follows an exponential distribution with unit mean. $\varrho_{n}$ represents shadow fading with a standard deviation of $\xi$. $\mathcal{A}$ is the path loss constant, $L_{k,t}^{n}$ is the distance between vehicle $n$ and BS, and $\iota$ is the decay exponent.
@@ III-C Computing model
where $T_{k,t}^{{\text{comp}},n}$ is the computing delay of vehicle $n$ to execute one iteration of local training, and $p_{k,t}^{{\text{comp}},n}$ is the CPU computing power of vehicle $n$, that can be calculated by Dynamic Voltage Frequency Scaling (DVFS) approach [52]-[54], i.e.,
@@ IV Optimization Problem
In the local training phase, higher CPU frequency allows for more iterations, which can enhance model performance but also increase energy consumption. In the uploading phase, vehicles upload the trained local model to BS. Higher transmission power reduces the upload time but increases energy consumption [57][58]. To save resources, we need to optimize the allocation of vehicle CPU frequencies and transmission power to minimize the overall energy consumption of all vehicles.
However, if the CPU frequency is too low, the time for a single local training iteration increases, resulting in fewer local iterations. Although this reduces energy consumption, it does not ensure the model’s performance [59][60]. Therefore, we also need to minimize the time for each training to ensure sufficient local iterations.
To deal with these two problems, our target is adjusted to minimize the total energy consumption of all vehicles and the max delay including model training and model transmission. The energy and delay for system performance can be calculated as:

### v2

@@ I Introduction
Internet of Vehicles (IoV) is a technology that connects vehicles, road and transportation facilities to Internet, enabling intelligent interaction and information sharing among them. Advances in IoV have made many practical applications possible, e.g., automatic navigation, meteorological information and road condition monitoring, which provide convenience to drivers and self-driving system, reduce the probability of accidents and improve the efficiency of the transportation system [1, 2, 3]. These applications need robust models, which require a lot of data for training. In IoV, vehicles can continuously capture the fresh image data along the moving road, including road conditions and environmental states through onboard sensors, where most of the data are captured in the form of images [4, 5, 6]. Moreover, vehicles also can store the image data locally and deal with them to realize recognition and classification, which provides essential information for self-driving and driver assistance systems to aid drivers in perceiving and understanding surroundings. For instance, a brand-new concept of underwater teleportation was proposed in [7], which can significantly improve the control stability and perceiving transparency of drivers via the human-on-the-loop system.
@@
However, the locally stored image data, denoted as local data, is usually private for vehicles, and vehicles are reluctant to share the data with each other to protect privacy. Since vehicles come from all directions, the local data they collect vary greatly. If the model, denoted as local model, is trained locally based on the local data, it may converge to a local optimum [8, 9]. In contrast to distributed training, Federated Learning (FL) performs a centered training and enables a Base Station (BS) to obtain a global model through aggregating local models from vehicles without sharing the local data, thereby providing a privacy-preserving solution and reducing the differences between local models [10, 11]. For local training, the traditional supervised learning requires the image data with labels [12, 13, 14], but data labeling requires significant manual effort. Especially in IoV, a large amount of unlabeled data is generated daily, further increasing the cost of manual labeling. The Self-Supervised Learning (SSL) leverages pretext to provide training based on the inherent properties of data, thus it can learn feature representations from the data itself without the need for manually annotated labels. The traditional SSL methods obtain a better training performance through generating and storing abundant negative samples, which occupies a lot of computing resource and storage space. Thus, traditional methods are not suitable for the resource and space limited IoV. The famous SSL method Momentum Contrast (MoCo) uses $k$ values to form a dictionary and stand for the reference of negative samples, which can save computing resource and storage space [15]. What’s more, MoCo uses dynamic dictionaries and momentum encoders to maintain a large and consistent set of negative samples, thereby improving the stability and quality of learned representations. However, employing MoCo in FSSL for local training will upload $k$ values to create a big queue in BS for aggregation, which may cause privacy leakage [16]. Moreover, in IoV, the high-velocity movement of vehicles often leads to motion blur. Specifically, excessive vehicle velocity may result in insufficient exposure time of the camera sensor, causing image blur [17, 18, 19]. Motion blur significantly reduces image clarity, making the features extracted from such images less accurate and leading to unstable feature representation during model training. At the same time, critical information in blurred images may be obscured, making it difficult for feature extraction algorithms to recognize and differentiate important image content, thus impacting the model’s accuracy and reliability. In FL, local models may perform poorly due to blur in the local data, thereby affecting the performance of the global model after aggregating. Data collected from different vehicles may exhibit significant distribution differences due to variations in environmental and motion conditions. This data heterogeneity usually leads lead to performance imbalance of the model across different vehicles, further exacerbating the challenges in model training and aggregation. To the best of our knowledge, there is no work considering the privacy leakage and motion blur for FSSL in IoV.
@@
The improved MoCo method Simplified Contrast (SimCo) uses dual temperature to control sample distribution to eliminate the usage of dictionary [20], thus it can greatly protect privacy and address the shortcomings of FedCo and MoCo in their integration with FL in IoV. Considering the negative impact of motion blur on model aggregation, we also propose a motion Blur resistant FSSL, refer to as BFSSL.
@@
It is noteworthy that increasing the Central Processing Unit (CPU) frequency of a vehicle results in more local iterations, which leads to higher accuracy of classification, while consuming more energy consumption. Higher transmission power of a vehicle reduces the transmission delay, while costing more energy. Simultaneously, low CPU frequency reduces the local iterations and the classification accuracy of the global model [21][22]. Additionally, low transmission power increases data interference, resulting in higher rates of transmission failure. Thus, it is critical to find a resource allocation scheme that allocates resources (i.e., CPU frequency and transmission power of vehicles) to minimize energy consumption and delay for the BFSSL, while increasing the success rate of transmission in IoV. Since the resource allocation problem is usually Non-convex in the complicated IoV, the traditional convex optimization methods are difficult to solve it. Deep Reinforcement Learning (DRL) is a good option to solve the Non-convex problem. In this paper, we propose the DRL-BFSSL algorithm, where vehicles employs SimCo to perform locally and sends local models to BS for blur-based aggregation. Meanwhile, BS allocates CPU frequency and transmission power to minimize the energy consumption and delay [^1] The source code has been released at: https://github.com/qiongwu86/DRL-BFSSL.
@@
  • In IoV, we utilize the level of motion blur as a crucial weighting factor on Independent and Identically Distributed (IID) and Non-IID datasets during model aggregation. Subsequently, aggregating models uploaded by training vehicles based on the blur level, we obtain a blur resistant global model.
  • We firstly formulate an optimization problem aiming to minimize the sum of energy consumption and delay. Subsequently, utilizing the Karush-Kuhn-Tucker (KKT) conditions to determine the optimal solution that jointly allocating bandwidth and CPU frequency. Finally, by the Soft Actor-Critic (SAC) algorithm, we identify the optimal allocation schedule for the CPU frequency and transmission power of each vehicle.
  • Given that low transmission power can result in data transmission failures, we use the data error rate to establish a minimum transmission power limit.
The rest of this paper is divided into several interrelated parts to comprehensively explore the discussed problem. In Section II, we will review prior work relevant to this study, providing background and context for subsequent discussions. In Section III, we describe the system model. In Section IV, we will formulate an optimization problem aiming to minimize the sum of energy consumption and delay and in Section V, we will introduce our proposed DRL-BFSSL algorithm. Section VI showcase and discuss the simulation results. Ultimately, we will summarize the paper and in Section VII.
@@ II-A Federated SSL
In [23], Jahan et al. proposed a FSSL which combines FL and the SSL method (i.e., Simple framework for Contrastive Learning of Representations (SimCLR)), to identify Mpox from skin lesion images [24]. In [25], Yan et al. proposed a FSSL which combines FL and the SSL method (i.e., Variational inference with adversarial learning for end-to-end text-to-speech (Vits)), and utilizes medical images stored in each hospital to train models. In [26], Xiao et al. proposed a FSSL which integrates Generative Adversarial Networks (GAN) as a pretext for SSL and combines it with FL to support automatic traffic analysis and synthesis over a large number of heterogeneous datasets. In [27], Li et al. proposed a FSSL which integrates You Only Look Once (YOLO) as a pretext for SSL and combines it with FL to process object detection in power operation sites. These FSSL methods mentioned above require a large number of samples for training, which imposes high demands on the computing and storage capabilities of devices. However, the computing and storage resources of vehicles are limited in IoV, thus these methods are not suitable for IoV [28].
@@ II-B DRL-based resource allocation scheme in FL
Some works have proposed DRL-based resource allocation schemes in IoV [33][34]. In [35], Zheng et al. proposed a resource allocation scheme based on DRL method to minimize the duration required for accomplishing tasks while respecting latency limitations in IoV. In [36], Hazarika et al. proposed a resource allocation scheme based on DRL method to allocate the power optimally in IoV. In [37], Guo et al. proposed a resource allocation scheme based on DRL method to minimize the cost of Road Side Unit (RSU) in IoV. However, the methods mentioned above have not considered the FL framework. There are also some works that consider the FL framework to address resource allocation problems based on DRL. In [38], He et al. combined FL and DRL method (i.e., Deep Deterministic Policy Gradient (DDPG)), to optimize both computation offloading and resource allocation. In [39], Zhang et al. combined FL and DRL method (i.e., DDPG), and proposed a dynamic computation offloading and resource allocation scheme to minimize the delay and energy consumption. In [40], Zhao et al. designed a multi-objective optimization problem to assist Hierarchical Federated Learning (HFL) with Mobile Edge Computing (MEC), aiming to achieve more accurate models and reduce overhead. In [41], Zhou et al. proposed a joint RL method based on local differential privacy. This method allows devices to add noise during local training, with the central server then designing the optimal resource allocation strategy. In [42], Zhang et al. formulated a complex non-convex problem with the goal of minimizing the latency and energy consumption required for task execution. They proposed a multimedia task offloading and resource allocation scheme based on Federated Deep Q-Network. The aforementioned methods highlight the crucial role of DRL algorithms in addressing complex non-convex problems. However, these algorithms are primarily designed for static environments. In scenarios characterized by high mobility and rapid environmental changes, these algorithms will not offer efficient solutions.
Overall, there is no work concerning the DRL-based resource allocation schemes for FSSL in IoV. Therefore, we further investigate a DRL-based resource allocation scheme to optimize the performance for the BFSSL in IoV, referred to as DRL-BFSSL.
@@ III System Model
In this section, we will introduce the system scenario and models. As shown in Fig. 1, we consider a urban scenario of a BS deployed at the center of the intersection and several vehicles driving in the coverage of the BS. When a vehicle arrives at the intersection, it turns left, turns right and drives straightly with probabilities $\varepsilon_{1}$, $\varepsilon_{2}$, $\varepsilon_{3}$, and $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}=1$, and then it keeps driving straightly. Vehicle velocity follows the truncated Gaussian distribution. Each vehicle is equipped with a camera to capture images during driving through the coverage area of the BS, and the vehicles coming from different directions capture different categories of images. $N$ training vehicles are randomly selected from the vehicles before crossing the intersection to ensure that vehicles have enough time for local training. Vehicle $n$, $n\in[1,N]$, is equipped with a camera and captures $Z$ image data with velocity $v_{0}^{n}$ before it enters the coverage area of the BS.
The entire process of DRL-BFSSL algorithm consists of $K^{\text{max}}$ episodes, each of which is divided into two task lines, one for BFSSL, where episode $k$ is divided into different rounds, and the other for DRL-based resource allocation, where episode $k$ is divided into different slots.
Specifically, as shown in Fig. 2, within each episode $k$, $k\in[1,K^{\text{max}}]$ , there are $S^{\text{max}}$ slots. Each green cell represents a slot, and each orange cell represents a round. One slot corresponds to one round. The difference is that at the end of each episode $k$, when entering the next episode $k+1$, slot $t$ is reset to slot 00 and start with slot $1$. In contrast, round $r$ is not reset and continues to accumulate until the whole process of DRL-BFSSL algorithm finished. Due to each episode containing $S^{\text{max}}$ slots, and round $r$ resets to 0 after the completion of the $K^{\text{max}}$ episodes, the relationship between slot $t$ in episode $k$ and round $r$ is as follows
@@
At the beginning of the DRL-BFSSL algorithm, BS stores the parameters of the global model, while each vehicle maintains a local model with the same structure as the global model. Additionally, each vehicle keeps images captured before training, referred to as local data. Vehicles also upload their velocities to the BS, and then BS estimates their positions with the received velocities. Based on these estimates, BS quickly allocates transmission powers and CPU frequencies to vehicles. In each round of SSL, each vehicle first downloads the transmission power, CPU frequency, and the parameters of the global model from BS, and applies the parameters to its local model. The vehicles then train their local models using the local data according to the downloaded CPU frequency. After training is complete, the vehicles upload the trained parameters of their local models to BS with the assigned transmission power. During the transmission process, varying signals transmitted by other vehicles may cause interference, leading to data errors that prevent the BS from receiving the uploaded local models. Finally, BS aggregates the received local models to obtain a new global model based on the motion blur. The above process is repeated in different episode to get a converged global model.
@@ III-A Mobility model
As the velocity of each vehicle $n$ follows the truncated Gaussian distribution to reflect real vehicle mobility, the velocities of different vehicles are IID. Let $v_{k,t}^{n}$ be the velocity of vehicle $n$, and $v^{\text{min}}$ and $v^{\text{max}}$ be the minimum and maximum velocity of vehicles, respectively. Thus the probability density function of $v_{k,t}^{n}$ is calculated as [43]
@@ III-B Transmission model
During the transmission process, the energy consumption of vehicle $n$ can be calculated as
@@
where $c_{k,t}^{n}$ is the transmission rate of vehicle $n$, which can be calculated according to Shannon’s theorem [44], i.e.,
@@
where $g_{n}$ represents the small-scale fast fading power, which follows an exponential distribution with unit mean. $\varrho_{n}$ represents shadow fading with a standard deviation of $\xi$. $\mathcal{A}$ is the path loss constant, $L_{k,t}^{n}$ is the distance between vehicle $n$ and BS, and $\iota$ is the decay exponent [46].
@@ III-C Computing model
where $T_{k,t}^{{\text{comp}},n}$ is the computing delay of vehicle $n$ to execute one iteration of local training, and $p_{k,t}^{{\text{comp}},n}$ is the CPU computing power of vehicle $n$, that can be calculated by Dynamic Voltage Frequency Scaling (DVFS) approach [47], i.e.,
@@
To evaluate the utility of CPU frequency, we use the inverse of energy consumption as the criterion for the utility function. The utility $U$ can be expressed as [48]
$U=\frac{1}{E_{k,t}^{{\text{comp}},n}}.$ (11)
@@ IV Optimization Problem
In the local training phase, higher CPU frequency allows for more iterations, which can enhance model performance but also increase energy consumption. In the uploading phase, vehicles upload the trained local model to BS. Higher transmission power reduces the upload time but increases energy consumption. To save resources, we need to optimize the allocation of vehicle CPU frequencies and transmission power to minimize the overall energy consumption of all vehicles [51].
However, if the CPU frequency is too low, the time for a single local training iteration increases, resulting in fewer local iterations. Although this reduces energy consumption, it does not ensure the model’s performance. Therefore, we also need to minimize the time for each training to ensure sufficient local iterations [52][53].
To deal with these two problems, our target is adjusted to minimize the total energy consumption of all vehicles and the max delay including model training and model transmission [54][55]. The energy and delay for system performance can be calculated as: